Accessibility Assistance

Skip to Content




Information Technology
George A. Smathers Libraries
University of Florida
Suite 1250
2124 NE Waldo Rd
Gainesville, FL 32609

Telephone: (352) 273-2800
Fax: (352) 392-2354

Information Technology Staff Meeting Notes
June 14, 2010

Present: Shruti Bagri, Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Matthew Daley, Debra Fetzer, Winston Harris, Michael Jay, Tom Minton, Chris Nicolich, Amy Polk, Patricia Ruwell, Josh Spurgin, Evan Wack
Announcements
  1. Smathers south tower entrance/exit door is working!
  2. President Machen announced merit raises: 4% faculty, 3% staff. Retroactive as of July 1st.
  3. Convocation input requested. Please send a list of any departmental accomplishments to Will. Deadline is July 1st.
  4. TSS Division meeting feedback - smartphones, more training. We will include these topics in the IT Update meeting this month (probably June 30).
Ongoing IT Updates

Orders
  1. Ordering will resume in mid-July.
Relocation
  1. Eastside - paint job is very poor. Peter (Facilities) sent them a list of items to be revisited.
  2. Wrong numbers on the offices? Trying to verify the accuracy. This will cause a problem with Telecom.
  3. IP ranges requested from CNS.
  4. Move date: July 12.
Hardware
  1. 60% of our furniture order has been delivered to ALF.
  2. EPO - not much to report. Will discuss at the IT Update meeting.
  3. Public printing - met with Jeff Smith. He has a good relationship with the Pharos folks. They have tapped him to resolve our issues. We have scheduled a meeting between us, Canon, Purchasing, and Aramark to resolve the issues. Stacey Ewing will conduct a meeting to get input from the Librarians,
  4. Videoconferencing - microphones picking up the loud AC - PPD engineers to look at the AC this week. Air vents are out of balance.
  5. Altiris - we have a good road map. Image map is working. This is a work in progress. We won't be at 100% for about a year.
  6. There is a bug with the security software in the public image. We contacted Clean Slate, they are sending a fix.
  7. Stacey Ewing - July 19 - training on time-limited software.
  8. Win7 - redid staff image. Housekeeping tasks added for Altiris.
  9. Group policy - working with group policy allows us to put permissions down to OCLC and ALEPH.
  10. Privilege authority - will be going out with the image. It is done and working fine.
  11. Evan is working on a general IT info document to push to all staff machines.
Software
  1. Two new projects: Open Access Publishing Fund form for Isabel Silver. It has a lot of odd requirements: the form has to be signed, all data has to go into a database. High priority; deadline: July 1st. (Chris Nicolich)
    Second new project is the Document Disposition database - to assist selected libraries in the SE region get rid of objects and let others claim it if they desire. (Shruti & Tom Bielicke)
  2. Database migration - going well. Lots of small jobs moved; some more complex ones as well. Grover & ITS have been moved. This is now on hold due to higher priority projects.
  3. Acquisitions macro suite ongoing. Michael has been working on a macro with Hang out at ALF. The machines out there have a very slow connection.
  4. ARL Stats - Tom is helping Steve as he needs it. Waiting on Steve at this point.
  5. HR project - launched last week, It is live now. Finished, until they want a report.
  6. OLE - position has been open for a week. Not too many applicants. Closes Friday.
Network/Server Administration
  1. Sharepoint - should be hearing about access this week.
  2. DHCP should be live everywhere, except East and ALF. ALF will need a private IP range.
  3. CNS virtualization - just a few left. Printing is not done yet. Will probably do it during break week (next week). Web server - looking for a tool to do the virtual directory migration. Hope to start this week. At some point, we will need to block folks from updating Web pages for 1-3 days. App server - will clone it, and then move it. Vivo servers - 3 of those to move. EPO? Unsure how to do it. ARC GIS server - migration? Dongle issue. Can upgrade to an USB dongle. Printer migration - PS2 to move to CNS management server.
  4. Streaming video - vendors can host.
ILS
  1. Meeting with FCLA this afternoon.
DLC General Issues
  1. OCR boxes not done yet.
  2. JPG2K8 issue resolved. JPG2K3 server used instead.
  3. Last step: Greenstone boxes today and tomorrow.
  4. Snapshots of the virtual server - hard to get permission to get them. Logan to talk to CNS about it.
Web Projects
  1. Open Access PDF is now a programming project.
  2. Regional Depository is now a programming project.
  3. Proxy server - Jason & Cliff Richmond are passing URLs through. What about setting up a second proxy server? Would need approval from Public Services folks. There is a ticket in Logan's queue for a test proxy box. Logan will update the ticket and get everyone involved.
  4. Working on a second theme for the WordPress site, per Rachel's suggestion.
  5. Sharepoint - waiting until it rolls.
  6. Web migration - Matthew is going through a 27-page list. We've asked some staff to move these pages in the past. Let them know the deadline one last time and if they still are not in compliance, let Will know. All links will have to be checked.